OpenStudy (anonymous):

find the reminder when f(x)=-2x^3+3x^3-x+1/2 is divided by 2x+1

OpenStudy (yttrium):

It's all bout remainder theorem. Do you know something about it?

OpenStudy (anonymous):

r u sure it is 3x^3??

OpenStudy (anonymous):

ot it is 3x^2?

OpenStudy (anonymous):

oopzz.....sorry itz 3x^2...

OpenStudy (yttrium):

@prasheelasingh , something about remainder theorem. Remember?

OpenStudy (anonymous):

is the remainder 2?

OpenStudy (yttrium):

maybe for some sort of reference, read this one http://www.purplemath.com/modules/remaindr.htm

OpenStudy (anonymous):

have u ever practiced dis type of division?

OpenStudy (anonymous):

itz quite difficult for me...

OpenStudy (yttrium):

Anyway, to solve this we must solve variable x by our (divisor I think: just take a look at 2x+1) 2x+1 = 0 Therefore, x = -1/2 After that, substitute the value we got into the polynomial equation (take a look at -2x^3+3x^2-x+1/2) Hence, \[-2(\frac{ -1 }{ 2 })^3 + 3(\frac{ -1 }{ 2 })^2 - \frac{ -1 }{ 2 } + \frac{ 1 }{ 2 }\] Then, \[-2(\frac{ -1 }{ 8 }) + 3(\frac{ 1 }{ 4 }) + 1\] \[\frac{ 1 }{ 4 }+\frac{ 3 }{ 4 }+1\] Therefore, the remainder is 2.

OpenStudy (yttrium):

The way I used is based from remainder theorem.

OpenStudy (anonymous):

i have tried it by.....evaluating -1/2 in the equation......... -2(-1/2)^3+3(-1/2)^2-(-1/2)+1/2=1.5

OpenStudy (anonymous):

opppzzz....... i got it.....

OpenStudy (yttrium):

May be just do some sort of rechecking. Or you can post your solution so we can both verify.

OpenStudy (anonymous):

finally............thanks....i was wrong......

OpenStudy (anonymous):

the reminder is = to 2.......

OpenStudy (yttrium):

No problem. :))